# How Does A Text Preprocessing Pipeline Affect Syntactic Ontology Matching?

### v6

@@ Abstract.
The classic text preprocessing pipeline, comprising Tokenisation, Normalisation, Stop Words Removal, and Stemming/Lemmatisation, has been implemented in many systems for syntactic ontology matching (OM). However, the lack of standardisation in text preprocessing creates diversity in mapping results. In this paper, we investigate the effect of the text preprocessing pipeline on syntactic OM in 8 Ontology Alignment Evaluation Initiative (OAEI) tracks with 49 distinct alignments. We find that Phase 1 text preprocessing (Tokenisation and Normalisation) is more effective than Phase 2 text preprocessing (Stop Words Removal and Stemming/Lemmatisation). We propose two novel approaches to repair unwanted false mappings caused by Phase 2 text preprocessing. One is an ad hoc logic-based repair approach that employs an ontology-specific check to find common words that cause false mappings. These words are stored in a reserved word set and applied before the text preprocessing. By leveraging the power of large language models (LLMs), we also propose a post hoc LLM-based repair approach. This approach utilises the strong background knowledge provided by LLMs to repair non-existent and counter-intuitive false mappings after the text preprocessing. It also overcomes the tendency towards unstable true mappings by injecting the classical text preprocessing pipeline via function calling. The experimental results show that these two approaches can improve the matching correctness and the overall matching performance.
@@ 1. Introduction
Ontology matching (OM) is essential for enabling interoperability between heterogeneous ontologies. An OM process usually takes two ontologies as input, discovers mappings between entities, and produces a set of correspondences (Euzenat and Shvaiko, 2013). A classical OM system usually contains syntactic matching, lexical matching, and semantic matching. This multilayer architecture has been implemented in several successful OM systems, such as LogMap (Jiménez-Ruiz and Cuenca Grau, 2011; Jiménez-Ruiz et al., 2011), AgreementMakerLight (AML) (Faria et al., 2013, 2014), and FCA-Map (Zhao et al., 2018; Li et al., 2021).
While syntactic matching captures “anchor mappings” and provides the foundation for the latter lexical and semantic matching, there are many strategies for extracting syntactic information from an ontology entity. These include older approaches Bag-of-Words (e.g. TF-IDF (Sammut and Webb, 2010)), popular word embedding models (e.g. Word2Vec (Mikolov et al., 2013)), and state-of-the-art language models (e.g. BERT (Devlin et al., 2019)). Despite the diversity of the models used, they all apply the text preprocessing pipeline to clean the text data before using the model.
The classical text preprocessing pipeline consists of a set of steps to segment, reconstruct, analyse, and process the information in the text, namely Tokenisation (T), Normalisation (N), Stop Words Removal (R), and Stemming/Lemmatisation (S/L) (Manning et al., 2008). T is the process of breaking the text into the smallest units (i.e. tokens). N is the process of transforming these different tokens into a single canonical form. R is the process of removing filler words that usually carry little meaning and can be omitted in most cases. S/L is used to deal with the grammatical variation of words, applying rules to find the simplest common form of the word. Table 1 shows examples of using the text preprocessing pipeline to find mappings in syntactic OM. In the examples, we use whitespace to split the tokens, normalise all the words into lowercase, remove the common stop words “has”, “a” and “is”, and apply stemming/lemmatisation to find the simplest word form “review” from “reviews” and “reviewing”.
@@
Table 1. Examples of using the classical text preprocessing pipeline in syntactic ontology matching: Tokenisation (T), Normalisation (N), Stop Words Removal (R), and Stemming/Lemmatisation (S/L).
A number of OM systems use the text preprocessing pipeline for syntactic OM, but few studies question its role in optimising the matching performance. Our study fills this gap. We first conduct a comprehensive experimental analysis of the text preprocessing pipeline used in syntactic OM across a wide range of domains, aiming to explain the behaviour of the text preprocessing pipeline. In each step, text preprocessing methods are evaluated for correctness and completeness. While several steps of the text preprocessing pipeline are currently less effective, we propose two novel pipeline repair approaches. These two methods utilise the logic of ontological representations and the power of large language models (LLMs), showing promising results for repairing the less effective text preprocessing pipeline. Specifically, this paper makes the following contributions:
  • We compare the performance of the text preprocessing pipeline on 8 Ontology Alignment Evaluation Initiative (OAEI) (OAEI Community, [n. d.]) datasets with 49 distinct alignments. While there is a significant improvement using Tokenisation (T) and Normalisation (N) in syntactic OM, Stop Words Removal (R) and Stemming/Lemmatisation (S/L) are currently less effective. We find common inappropriate stop words removal, over-stemming, and over-lemmatisation.
@@ 2. Related Work
Syntactic OM considers only the meaning of an entity’s name or label, ignoring its lexical and structural context in an ontology. Correct syntactic matches are often implicit and usually require extra human observation and domain knowledge (Euzenat and Shvaiko, 2013). Text preprocessing is introduced to automate this process. The use of text preprocessing in syntactic OM can be traced back to the early stages of OM systems, which were initially developed and widely used as a basic component to generate linguistic-based mappings. AgreementMaker (Cruz et al., 2009) has a normaliser to unify the textual information of the entities. SAMBO (Lambrix and Tan, 2006) uses the Porter Stemmer for each word to improve the similarity measure for terms with different prefixes and suffixes. In RiMOM (Li et al., 2009), the context information of each entity is viewed as a document. The text in each document is preprocessed with tokenisation, stop words removal, and stemming.
Recently, machine learning (ML) models have emerged for modern OM systems. While text preprocessing remains useful, its role is more focused on normalising the text that becomes the input to the model. For example, BERTMap (He et al., 2022) uses BERT’s inherent WordPiece tokeniser to build the subword of each entity. DeepOnto (He et al., 2024) extends normalisation to axioms using EL embedding models (Kulmanov et al., 2019). The ML extension of LogMap (Chen et al., 2021) reuses the seed mappings of traditional systems, where each entity is split into its component words and the mapping is based on pairwise component similarities. However, to the best of our knowledge, most of the literature implements a preprocessing method without explaining why a specific method is chosen, and no studies have evaluated the effect of the entire text preprocessing pipeline on syntactic OM.
@@ 3. Preliminaries
Given a source ontology ($O_{s}$) and a target ontology ($O_{t}$), OM establishes mappings between pairs of entities drawn from each of the two ontologies. A correspondence (a.k.a. a mapping) is defined as a 4-tuple $(e_{1},e_{2},r,c)$, where $e_{1}\in O_{s}$ and $e_{2}\in O_{t}$. $r$ is the relationship between two matched entities $e_{1}$ and $e_{2}$, and $c\in[0,1]$ is the confidence of the correspondence (or may be absent). The relationship $r$ in OM tasks can be equivalence ($\equiv$), subsumption ($\subseteq$), or other more complex relationships. In this study, we focus only on the equivalence relationship ($\equiv$) and evaluate the effect of text preprocessing on syntactic matching to produce the “anchor mappings” on which to base any subsequent lexical and semantic matching. We address only equivalence because subsumption is typically dealt with in a later semantic and structural matching phase of OM. An alignment (A) is a set of candidate correspondences, while a reference (R) is a set of gold-standard correspondences verified by domain experts (i.e. the ground truth alignment).
@@ 4.1. Experiment Setup
Figure 1 shows the experiment setup to analyse the text preprocessing pipeline. Alignment (A) is generated via the text preprocessing pipeline. For both $O_{s}$ and $O_{t}$, we retrieve the entities from classes (owl:Class) and properties (both owl:ObjectProperty and owl:DatatypeProperty). For those ontologies where the names of the concepts are not textual (e.g. a code or numerical identifier), we retrieve the meaningful text from entity labels (rdfs:label or other annotation properties) instead. Then, we apply the text preprocessing pipeline method $f(\cdot)$ on each entity $e_{1}\in O_{s}$ and $e_{2}\in O_{t}$. If $f(e_{1})=f(e_{2})$, we store the correspondence in the corresponding alignment file. We compare the generated alignment (A) with the reference (R) to evaluate the performance of the text preprocessing pipeline on syntactic OM.
@@
    Selected OAEI tracks: We select 8 OAEI tracks stored in the Ontology Matching Evaluation Toolkit (MELT) (Hertling et al., 2019) public repository (retrieved January 1, 2025). Table 2 shows the details of the selected tracks. Each track may contain more than one alignment corresponding to different pairs of ontologies. For example, the Largebio Track has multiple reference files, pairing FMA (Rosse and Mejino, 2003) and NCI (Golbeck et al., 2003), FMA and SNOMED (Donnelly, 2006), and SNOMED and NCI, respectively. The number of alignments for each track is given in the table. There are 49 distinct alignments evaluated in this study.
@@
Compound words are frequently used in ontology naming conventions. For example, compound words such as “art gallery” can be formatted as the Pascal case “ArtGallery” used in YAGO (Suchanek et al., 2007) or the Snake case “art_gallery” used in Wikidata (Vrandečić and Krötzsch, 2014). Figure 3 shows the proportion of compound words in the selected OAEI tracks. For comparisons between entities with compound words, one approach is to use their tokens, where “XY” is equivalent to “YX” because they share the same concepts X and Y. However, our comparison considers both concepts and their order, so that “XY” is not equivalent to “YX”. Although concurrent “XY” and “YX” are rare in the matching process, we assume that it could happen and so we take account of the order of concepts in compound word names to ensure a fair comparison in our experiments.
Figure 3. Proportion of compound words. We exclude entities with more than 15 number of words because their proportion is less than 1%.
@@ 5.3. Ad Hoc Logic-Based Repair
The ad hoc logic-based repair applies before the text preprocessing. One critical step in our approach is to assemble a set of reserved words that may cause FPs in text preprocessing, and to exclude these words from text preprocessing. The selection criteria are based on two widely agreed assumptions: (1) there are no duplicate entities within a single ontology; and (2) ontologies that represent the same domain knowledge tend to use similar terminologies (OBO Foundry, [n. d.]). This approach is context-based because pairs of words may have the same or different meanings in different contexts. Based on these two assumptions, we propose a simple Algorithm 1 to assemble the reserved word set. For both $O_{s}$ and $O_{t}$, the algorithm first iterates on all pairs of entities $e_{i},e_{j}$ in the ontology, respectively. For a specific text preprocessing method $f(\cdot)$, if $f(e_{i})=f(e_{j})$, we retrieve the different words between $e_{i}$ and $e_{j}$ and store them in the reserved word set. If a word appears in the reserved set, the text preprocessing pipeline skips the operation for this word. To simplify the reserved word set, we remove the words where $f(w)=w$ because their occurrence in the reserved word set has no effect on mapping results.
@@
0:  Source Ontology $O_{s}$, Target Ontology $O_{t}$, Text Preprocessing Pipeline Method $f(x)$
0:  $Reserved\_Word\_Set$
@@
        $Reserved\_Word\_Set\leftarrow differ(e_{i},e_{j})$
@@
  /* Phase 2: Find immutable words in $Reserved\_Word\_Set$/
  for $Word\ w\in Reserved\_Word\_Set$ do
@@
  return  $Reserved\_Word\_Set$
@@ 5.4. Post Hoc LLM-Based Repair
    Figure 9. Discovery rate of TPs using LLMs on the Largebio Track.
    Figure 10. The discovery rate of FPs using LLMs on the Largebio Track.
@@
    We compare LLMs with the classical text preprocessing pipeline in detecting TPs and FPs, which are two main contributors to precision and recall in matching performance. The true mappings (TPs) are retrieved from the reference (R) where the mappings can be found by syntactic OM. The false mappings (FPs) are retrieved from the alignment (A) generated by the classic text preprocessing pipeline but exclude TPs above.
@@
Figure 9 compares LLMs with the classical text preprocessing pipeline in detecting TPs. LLMs and the text preprocessing pipeline are expected to answer “yes” for TPs. The discovery rate is defined as the number of “yes” answers compared to the total number of TPs. For LLMs, we directly prompt LLMs to determine whether two entities are identical or not. For the text preprocessing pipeline, two entities are considered to be matched if they have the same results after the text preprocessing. Here we use only Tokenisation (T) and Normalisation (N) to preprocess the text because Stop Words Removal (R) and Stemming/Lemmatisation (S/L) are shown to be less effective in our analysis in Section 4. Although some API-accessed commercial LLMs have shown the capability to detect TPs without a text preprocessing step, this behaviour varies across different LLMs and prompt templates. For different LLMs, open-source LLMs perform better than commercial API-accessed LLMs. It could be problematic for some commercial API-accessed LLMs (e.g. claude-3-haiku) to find two identical entities if text preprocessing is not applied prior to the comparison. They may consider entities with different formatting that are used in different contexts and therefore have different meanings. For different prompt templates, few-shot (PT2) significantly improves correctness, but self-reflection (PT3) may even hamper performance. The combination of few-shot and self-reflection (PT4) may even perform worse than the simple basic prompt (PT1). We believe that this phenomenon is mainly attributed to “overthinking”, a type of LLM hallucination in which LLMs could over-analyse user input and generate unnecessarily complex details (Chen et al., 2025). We also need to note here that the example generated in the few-shot prompt is based on our observation of the existing TPs. Finding a good example in real-world applications is not easy because ontology entities can have different formatting by nature. On the other hand, using the classic text preprocessing pipeline via function calling could help LLMs produce a stable result with a high discovery rate (almost 100% from our observation) across different models and prompt templates.
@@
Figure 10 compares LLMs with the classical text preprocessing pipeline in detecting FPs. LLMs and the text preprocessing pipeline are expected to answer “no” for FPs. The discovery rate is defined as the number of “no” answers compared to the total number of FPs. Specifically, these FPs evaluated are non-trivial mappings that cannot be detected by the classic text preprocessing method. Here we use LLMs to detect FPs generated by the Lancaster Stemmer (SL). SL has shown a lower precision compared to other text preprocessing methods, which is caused by the method producing a significant number of FPs. LLMs have shown a strong capability to discover the FPs produced by the classic text preprocessing pipeline. For different models, API-accessed commercial LLMs are better than open-source LLMs and reach a higher discovery rate of more than 80%. For different prompt templates, both few-shot (PT2), self-reflection (PT3), and a combination of the above (PT4) are better than a simple basic prompt (PT4). While mapping repair is usually non-trivial and requires domain expertise, LLMs have shown potential for performing a high-quality post hoc correction without human intervention.
@@
For syntactic OM, we can see that LLMs are not good at detecting true mappings, but they are good at detecting false mappings. Therefore, we propose our generic post hoc LLM-based repair with (1) employing the classical text preprocessing pipeline via function calling (OpenAI, 2023) (a.k.a. tool calling) to detect TPs. This means that the LLMs act as an interface, and the mapping judgement is based on the programme logic defined in the text preprocessing method. This provides stable results for text preprocessing when finding FPs. LLM prompting is not suitable for this task, even with complex prompting strategies. (2) Using LLM prompting to detect FPs. The prompt template can be simple like “Is A equivalent to B?”, as we cannot see significant improvement using complex prompting strategies (except Meta Llama models).
5.5. Evaluation & Ablation Study

### v7

@@ Abstract.
The classic text preprocessing pipeline, comprising Tokenisation, Normalisation, Stop Words Removal, and Stemming/Lemmatisation, has been implemented in many systems for syntactic ontology matching (OM). However, the lack of standardisation in text preprocessing creates diversity in mapping results. In this paper, we investigate the effect of the text preprocessing pipeline on syntactic OM in 8 Ontology Alignment Evaluation Initiative (OAEI) tracks with 49 distinct alignments. We find that Phase 1 text preprocessing (Tokenisation and Normalisation) is more effective than Phase 2 text preprocessing (Stop Words Removal and Stemming/Lemmatisation). We propose two novel approaches to repair unwanted false mappings caused by Phase 2 text preprocessing. One is an ad hoc logic-based repair approach that employs an ontology-specific check to find common words that cause false mappings. These words are stored in a reserved word set and applied before the text preprocessing. By leveraging the power of large language models (LLMs), we also propose a post hoc LLM-based repair approach. This approach utilises the strong background knowledge provided by LLMs to repair non-existent and counter-intuitive false mappings after the text preprocessing. It also overcomes the tendency towards unstable true mappings by injecting the classic text preprocessing pipeline via function calling. The experimental results show that these two approaches can improve the matching correctness and the overall matching performance.
@@ 1. Introduction
Ontology matching (OM) is essential for enabling interoperability between heterogeneous ontologies. An OM process usually takes two ontologies as input, discovers mappings between entities, and produces a set of correspondences (Euzenat and Shvaiko, 2013). A classic OM system usually contains syntactic matching, lexical matching, and semantic matching. This multilayer architecture has been implemented in several successful OM systems, such as LogMap (Jiménez-Ruiz and Cuenca Grau, 2011; Jiménez-Ruiz et al., 2011), AgreementMakerLight (AML) (Faria et al., 2013, 2014), and FCA-Map (Zhao et al., 2018; Li et al., 2021).
Syntactic matching captures “anchor mappings” and provides the foundation for the latter lexical and semantic matching. There are many strategies for extracting syntactic information from an ontology entity. These include older approaches Bag-of-Words (e.g. TF-IDF (Sammut and Webb, 2010)), popular word embedding models (e.g. Word2Vec (Mikolov et al., 2013)), and state-of-the-art language models (e.g. BERT (Devlin et al., 2019)). Despite the diversity of the strategies used, they all apply the text preprocessing pipeline to clean the text data before using the model.
The classic text preprocessing pipeline consists of a set of steps to segment, reconstruct, analyse, and process the information in the text, namely Tokenisation (T), Normalisation (N), Stop Words Removal (R), and Stemming/Lemmatisation (S/L) (Manning et al., 2008). T is the process of breaking the text into the smallest units (i.e. tokens). N is the process of transforming these different tokens into a single canonical form. R is the process of removing filler words that usually carry little meaning and can be omitted in most cases. S/L is used to deal with the grammatical variation of words, applying rules to find the simplest common form of the word. Table 1 shows examples of using the text preprocessing pipeline to find mappings in syntactic OM. In the examples, we use whitespace to split the tokens, normalise all the words into lowercase, remove the common stop words “has”, “a” and “is”, and apply stemming/lemmatisation to find the simplest word form “review” from “reviews” and “reviewing”.
@@
Table 1. Examples of using the classic text preprocessing pipeline in syntactic ontology matching: Tokenisation (T), Normalisation (N), Stop Words Removal (R), and Stemming/Lemmatisation (S/L).
A number of OM systems use the text preprocessing pipeline for syntactic OM, but few studies question its role in optimising the matching performance. Our study fills this gap. We first conduct a comprehensive experimental analysis of the text preprocessing pipeline used in syntactic OM across a wide range of domains, aiming to explain the behaviour of the text preprocessing pipeline. In each step, text preprocessing methods are evaluated for correctness and completeness. While several steps of the text preprocessing pipeline are currently less effective, we propose two novel pipeline repair approaches. These two methods utilise the logic of ontological representations and the power of large language models (LLMs), showing promising results for repairing weaknesses of the text preprocessing pipeline. Specifically, this paper makes the following contributions:
  • We compare the performance of the text preprocessing pipeline on 8 datasets of the Ontology Alignment Evaluation Initiative (OAEI) (OAEI Community, [n. d.]) with 49 distinct alignments. While there is a significant improvement using Tokenisation (T) and Normalisation (N) in syntactic OM, Stop Words Removal (R) and Stemming/Lemmatisation (S/L) are currently less effective. We find common inappropriate stop words removal, over-stemming, and over-lemmatisation.
@@ 2. Related Work
Syntactic OM considers only the meaning of an entity’s name or label, ignoring its lexical and structural context in an ontology. Correct syntactic matches are often implicit and usually require extra human observation and domain knowledge (Euzenat and Shvaiko, 2013). Text preprocessing is introduced to automate this process. The use of text preprocessing in syntactic OM can be traced back to the early stages of OM systems, which were initially developed and widely used as a basic component to generate linguistic-based mappings. AgreementMaker (Cruz et al., 2009) has a normaliser to unify the textual information of the entities. SAMBO (Lambrix and Tan, 2006) uses the Porter Stemmer for each word to improve the similarity measure for terms with different prefixes and suffixes. In RiMOM (Li et al., 2009), the context information of each entity is viewed as a document. The text in each document is preprocessed with tokenisation, stop words removal, and stemming. Machine learning (ML) models have emerged for modern OM systems. While text preprocessing remains useful, its role is more focused on normalising the text that becomes the input to the model. For example, BERTMap (He et al., 2022) uses BERT’s inherent WordPiece tokeniser to build the subword of each entity. DeepOnto (He et al., 2024) extends normalisation to axioms using EL embedding models (Kulmanov et al., 2019). The ML extension of LogMap (Chen et al., 2021) reuses the seed mappings of traditional systems, where each entity is split into its component words and the mapping is based on pairwise component similarities. However, to the best of our knowledge, most of the literature implements a preprocessing method without explaining why a specific method is chosen, and no studies have evaluated the effect of the entire text preprocessing pipeline on syntactic OM.
Early explorations in (He et al., 2023; Norouzi et al., 2023) show the potential of using LLMs for OM, but matching by LLM prompting alone commonly results in poor performance. OLaLa (Hertling and Paulheim, 2023) proposes a prototype and states that the LLMs for OM include a number of decisions on model and parameter combinations, and there is no one-model-fits-all solution. Works on both Agent-OM (Qiang et al., 2023) and LLM4OM (Giglou et al., 2024) point out that LLMs have limited context length and inherited hallucinations. While LLM4OM highlights the importance of integrating the Retrieval-Augmented Generation (RAG) (Lewis et al., 2020), Agent-OM introduces an agent-based framework and enhances the compatibility of LLMs via function calling (OpenAI, 2023) for OM to improve matching performance. The authors in (Taboada et al., 2025) state that the use of LLMs for OM has a high computational overhead and apply a prioritised depth-first search (PDFS) strategy on the retrieve-then-prompt pipeline to optimise candidate selection. Recent works in (Amini et al., 2024; Zamazal, 2024; Sousa et al., 2025) also utilise LLMs and LLM embeddings for complex OM. However, these works focus on building a sophisticated framework for using LLMs for OM, and less attention has been paid to the subtask of using LLMs for syntactic matching. For syntactic OM, the challenges of using LLMs are twofold. On one hand, LLM hallucinations may produce incorrect “seed mappings”. Such mappings are toxic for the subsequent lexical and semantic matching. On the other hand, it is difficult to select a generic LLM prompting strategy from a number of LLM prompting templates and strategies.
@@ 3. Preliminaries
An ontology provides a formalised conceptualisation of domain knowledge (Gruber, 1993). The basic component of an ontology is the entity. Each entity is assigned a unique identifier (i.e. IRI) and represents a unique concept of the domain. Entities can be a concept (i.e. class), relation (i.e. property), or instance (i.e. individual) (Noy et al., 2001). Ideally, the entity names are not duplicated within a single ontology to minimise confusion and overlap. It is recommended that ontologies representing similar domain knowledge use similar terminologies for entity names (OBO Foundry, [n. d.]).
Ontology matching (OM) aims to capture and align similar concepts that occur in different ontologies. Given a source ontology ($O_{s}$) and a target ontology ($O_{t}$), OM establishes mappings between pairs of entities drawn from each of the two ontologies. A correspondence (a.k.a. a mapping) is defined as a 4-tuple $(e_{1},e_{2},r,c)$, where $e_{1}\in O_{s}$ and $e_{2}\in O_{t}$. $r$ is the relationship between two matched entities $e_{1}$ and $e_{2}$, and $c\in[0,1]$ is the confidence of the correspondence (or may be absent). The relationship $r$ in OM tasks can be equivalence ($\equiv$), subsumption ($\subseteq$), or other more complex relationships (Euzenat and Shvaiko, 2013). OM is a complex process that combines syntactic matching, lexical matching, and semantic matching. Each level of matching may use different matching techniques, for example, text preprocessing for syntactic matching, text vectorisation for lexical matching, and logical reasoning for semantic matching. This study concentrates on the common practice of using the text preprocessing pipeline in syntactic OM. We focus only on the equivalence relationship ($\equiv$) and evaluate the effect of text preprocessing on syntactic matching to produce the “anchor mappings” on which to base any subsequent lexical and semantic matching. We address only equivalence because subsumption is typically dealt with in a later semantic and structural matching phase of OM.
@@ 4.1. Experiment Setup
Figure 1 shows the experiment setup to analyse the text preprocessing pipeline. Alignment (A) is generated via the text preprocessing pipeline. For both $O_{s}$ and $O_{t}$, we retrieve the entities from classes (owl:Class) and properties (both owl:ObjectProperty and owl:DatatypeProperty). For those ontologies where the names of the concepts are not textual (e.g. a code or numerical identifier), we retrieve the meaningful text from entity labels (rdfs:label or other annotation properties) instead. Then, we apply the text preprocessing pipeline method $f(\cdot)$ on each entity $e_{1}\in O_{s}$ and $e_{2}\in O_{t}$. If $f(e_{1})=f(e_{2})$, we store the correspondence in the corresponding alignment file. While OM methods will often follow text preprocessing with a more sophisticated strategy to determine equivalence of entities, we define equivalence to mean identical text generated by preprocessing alone. This enables us to directly evaluate the preprocessing performance as a standalone contribution to OM. We compare the generated alignment (A) with the reference (R) to evaluate the performance of the text preprocessing pipeline on syntactic OM.
@@
[…]
Chart of experiment setup to analyse the text preprocessing pipeline.
@@
    Selected OAEI tracks: We select 8 OAEI tracks stored in the Ontology Matching Evaluation Toolkit (MELT) (Hertling et al., 2019) public repository (retrieved January 1, 2025). Table 2 shows the details of the selected tracks. Each track may contain more than one alignment corresponding to different pairs of ontologies. For example, the Largebio Track has multiple reference files, pairing FMA (Rosse and Mejino, 2003) and NCI (Golbeck et al., 2003), FMA and SNOMED (Donnelly, 2006), and SNOMED and NCI, respectively. The number of alignments for each track is given in the table. There are 49 distinct alignments evaluated in this study. Figure 2 shows the number of entities in each alignment.
@@
    Figure 2. Number of entities in each alignment.
[…]
    Chart of number of entities in each alignment.
    Compound words are frequently used in ontology naming conventions. For example, compound words such as “art gallery” can be formatted as the Pascal case “ArtGallery” used in YAGO (Suchanek et al., 2007) or the Snake case “art_gallery” used in Wikidata (Vrandečić and Krötzsch, 2014). Figure 3(b) shows the frequency distribution of compound words in the selected OAEI tracks. For comparisons between entities with compound words, one approach is to use their tokens, where “XY” is equivalent to “YX” because they share the same concepts X and Y. However, our comparison considers both concepts and their order, so that “XY” is not equivalent to “YX”. Although concurrent “XY” and “YX” are rare in the matching process, we assume that it could happen and so we take account of the order of concepts in compound word names to ensure a fair comparison in our experiments.
    Figure 3. Frequency distribution of compound words. We exclude entities with more than 15 compound words because their proportion is less than 1%.
[…]
    Chart of frequency distribution of compound words.
@@ 5.3. Ad Hoc Logic-Based Repair
The ad hoc logic-based repair applies before the text preprocessing. One critical step in our approach is to assemble a set of reserved words that may cause FPs in text preprocessing, and to exclude these words from text preprocessing. The selection criteria are based on two ontology features: (1) entity names (class or property) are not duplicated within a single ontology; and (2) ontologies that represent the same domain knowledge tend to use similar terminologies (see Section 3). This approach is context-based because pairs of words may have the same or different meanings in different contexts. Based on these two features, we propose a simple Algorithm 1 to assemble the reserved word set. For both $O_{s}$ and $O_{t}$, the algorithm first iterates on all pairs of entities $e_{i},e_{j}$ in the ontology, respectively. For a specific text preprocessing method $f(\cdot)$, if $f(e_{i})=f(e_{j})$, we retrieve the different words between $e_{i}$ and $e_{j}$ and store them in the reserved word set. If a word appears in the reserved set, the text preprocessing pipeline skips the operation for this word. To simplify the reserved word set, we remove the words where $f(w)=w$ because their occurrence in the reserved word set has no effect on mapping results.
@@
0:  Source Ontology $O_{s}$, Target Ontology $O_{t}$, Text Preprocessing Pipeline Method $f(\cdot)$
0:  Reserved_Word_Set $S$
@@
        $diff(e_{i},e_{j})=(e_{i}\setminus e_{j})\cup(e_{j}\setminus e_{i})$
        $S=S+diff(e_{i},e_{j})$
@@
  /* Phase 2: Find immutable words in $S$/
  for $Word\ w\in S$ do
@@
  return  Reserved_Word_Set $S$
@@ 5.4. Post Hoc LLM-Based Repair
    We compare LLMs with the classic text preprocessing pipeline in detecting TPs and FPs, which are two main contributors to precision and recall in matching performance. The true mappings (TPs) are retrieved from the reference (R) where the mappings can be found by syntactic OM. The false mappings (FPs) are retrieved from the alignment (A) generated by the classic text preprocessing pipeline but exclude TPs above.
@@
Figure 10(l) compares LLMs with the classic text preprocessing pipeline in detecting TPs. LLMs and the text preprocessing pipeline are expected to answer “yes” for TPs. The discovery rate is defined as the number of “yes” answers compared to the total number of TPs. For LLMs, we directly prompt LLMs to determine whether two entities are identical or not. For the text preprocessing pipeline, two entities are considered to be matched if they have the same results after the text preprocessing. Here we use only Tokenisation (T) and Normalisation (N) to preprocess the text because Stop Words Removal (R) and Stemming/Lemmatisation (S/L) are shown to be less effective in our analysis in Section 4. Although some API-accessed commercial LLMs have shown the capability to detect TPs without a text preprocessing step, this behaviour varies across different LLMs and prompt templates. For different LLMs, open-source LLMs perform better than commercial API-accessed LLMs. It could be problematic for some commercial API-accessed LLMs (e.g. claude-3-haiku) to find two identical entities if text preprocessing is not applied prior to the comparison. They may consider entities with different formatting that are used in different contexts and therefore have different meanings. For different prompt templates, few-shot (PT2) significantly improves correctness, but self-reflection (PT3) may even hamper performance. The combination of few-shot and self-reflection (PT4) may even perform worse than the simple basic prompt (PT1). We believe that this phenomenon is mainly attributed to “overthinking”, a type of LLM hallucination in which LLMs could over-analyse user input and generate unnecessarily complex details (Chen et al., 2025). We also need to note here that the example generated in the few-shot prompt is based on our observation of the existing TPs. Finding a good example in real-world applications is not easy because ontology entities can have different formatting by nature. On the other hand, using the classic text preprocessing pipeline via function calling could help LLMs produce a stable result with a high discovery rate (almost 100% from our observation) across different models and prompt templates.
Figure 10. Discovery rate of TPs using LLMs on the Largebio Track.
[…]
Discovery rate of TPs using LLMs on the Largebio Track.
Figure 11. The discovery rate of FPs using LLMs on the Largebio Track.
[…]
Discovery rate of FPs using LLMs on the Largebio Track.
@@
Figure 11(l) compares LLMs with the classic text preprocessing pipeline in detecting FPs. LLMs and the text preprocessing pipeline are expected to answer “no” for FPs. The discovery rate is defined as the number of “no” answers compared to the total number of FPs. Specifically, these FPs evaluated are non-trivial mappings that cannot be detected by the classic text preprocessing method. Here we use LLMs to detect FPs generated by the Lancaster Stemmer (SL). SL has shown a lower precision compared to other text preprocessing methods, which is caused by the method producing a significant number of FPs. LLMs have shown a strong capability to discover the FPs produced by the classic text preprocessing pipeline. For different models, API-accessed commercial LLMs are better than open-source LLMs and reach a higher discovery rate of more than 80%. For different prompt templates, both few-shot (PT2), self-reflection (PT3), and a combination of the above (PT4) are better than a simple basic prompt (PT4). While mapping repair is usually non-trivial and requires domain expertise, LLMs have shown potential for performing a high-quality post hoc correction without human intervention.
For syntactic OM, we can see that LLMs are not good at detecting true mappings, but they are good at detecting false mappings. Therefore, we propose our generic post hoc LLM-based repair with
  1. (1)
@@
    Employ the classic text preprocessing pipeline via function calling (a.k.a. tool calling) to detect TPs. This means that the LLMs act as an interface, and the mapping judgement is based on the programme logic defined in the text preprocessing method. This provides stable results for text preprocessing when finding FPs. LLM prompting is not suitable for this task, even with complex prompting strategies.
  2. (2)
    Use LLM prompting to detect FPs. The prompt template can be simple like “Is A equivalent to B?”, as we cannot see significant improvement using complex prompting strategies (except Meta Llama models).
5.5. Evaluation